Mathematics 242 – Principles of Analysis Definitions Quiz 9 – April 29, 2011

Your Name:

A) (10) Let f be a bounded function on [a, b] and given

\underline{∫} _{a}^{b}f = lub\{L_{P}(f) | P a partition of [a, b]\}

\overline{∫} _{a}^{b}f = glb\{U_{P}(f) | P a partition of [a, b]\}.

What does it mean to say that f is integrable on [a, b]?

B) (10) What does the Axiom of Completeness for R say?